# Failure Makes the Agent Stronger: Enhancing Accuracy through Structured Reflection for Reliable Tool Interactions

### v2

@@ 4.1 Experiment Settings
We conduct training on our self-constructed Tool-Reflection-Bench. After human supervision and post-editing, we retained approximately 5k samples in JSONL format to ensure compatibility with RL training under the Swift [34] framework.
@@
To verify the generalizability of Tool-Reflection-Bench and our training methodology, we conducted experiments using Llama3.1-8B [35], Qwen2.5-7B-Instruct [36], and Qwen3-4B [37] as base models.
Evaluation Metrics.
We evaluated multi-turn tool-calling performance using the Berkeley Function Calling Leaderboard (BFCL) v3 [17], with evaluation dimensions covering multi-turn-base, multi-turn-long-context, multi-turn-miss-func, and multi-turn-miss-param, and the evaluation metric being accuracy. To assess the model’s repair capability when tool calls fail, we used Tool-Reflection-Bench, with the evaluation metric being repair rate, Repair@n denotes that for the same data instance, if at least one out of n trials succeeds, the metric is recorded as 1; otherwise, it is 0.
4.2 Experiment Results
4.2.1 Result on BFCL v3
Comparison with base models.
We conduct performance evaluation on the multi-turn category of BFCL v3 to assess the benefits of enhancing the model’s self-reflection capability in multi-turn tool calling, the detailed results are showed in Table. 1. Compared the results against the corresponding base models. The most striking lift appears on Llama-3.1-8B: Base rises from 5.0 to 9.5 (+95%) and Long_Context from 4.5 to 7.0 (+56%). Qwen2.5-7B shows the largest Miss_Param gain (9.0 $\to$ 13.5, +50%), evidencing stronger parameter repair. Qwen3-4B attains an amazing absolute Multi-turn Overall (20.75, +28%) with a sizable Long_Context improvement (+48%). In contrast, its Miss_Func gain is modest (19.0 $\to$ 19.5, +3%), indicating tool selection remains comparatively harder—consistent with our method’s emphasis on reflection-driven parameter correction and long-context recovery.
4.2.2 Result on Tool-Reflection-Bench
As shown in Table. 2, across open-source baselines, repair rates are low at one try (Repair@1 $\leq$ 9.6%) and only mildly improve with more tries. Training with our method yields consistent gains for all bases: Llama-3.1-8B-Instruct jumps from $0.7/5.1/6.8$ to 4.7/20.5/26.4 (Repair@1/3/5), a large improvement especially at higher $n$; Qwen2.5-7B-Instruct improves from $2.4/6.1/8.0$ to 9.3/10.3/11.4; Qwen3-4B-Instruct rises from $9.6/10.6/10.6$ to 14.9/18.5/19.5 (best Repair@1 among our models). All finetuned models surpass the closed-source LongCat-Lite-8K-Chat across $n\!\in\!\{1,3,5\}$, indicating that our reflection-aware reward and RL objective substantially enhance repairability and yield more reliable multi-try recovery. It is also worth noting that when tool calls fail and require repair, our method achieves superior performance compared to closed-sourced models of the same scale such as LongCat-Lite-8K-Chat [38], GPT-4o-mini [39, 40], GPT-4.1-mini [41].
@@ 5 Conclusion
6 Prompt for Perturbation-based Disruptions
@@
6.1 Prompt for Call-Order Swap
@@
6.2 Prompt for Redundant Call
@@
6.3 Prompt for Missing Call
@@
6.4 Prompt for Argument Error
@@
7 Training Data Case Study
In this section, we extract one complete sample from each of the four perturbation modes for analysis.
7.1 Case Study of Call-Order Swap
@@
7.2 Case Study of Redundant Call
@@
7.3 Case Study of Missing Call
@@ 7.4 Case Study of Argument Error
The tool schema exposes multiple functions with schema–strict parameters (e.g., check_plant_water_level(plant_location:string), start_watering(plant_location:string, duration:number), start_trimming(hedge_location:string), …). The user requests two primary actions in the backyard: (i) trim hedges and (ii) water all potted plants for about 10 minutes; afterwards ensure plants have enough water and dispose clippings.
@@
  • start_watering(plant_location="backyard", duration=10)
  • start_trimming(hedge_location="backyard")
@@
8 Test Data Case Study
In this section, we present two representative test cases and their corresponding evaluation results as a case study, providing an intuitive demonstration of the effectiveness of our method and the model’s self-reflection capability for tool-call repair.
8.1 Case I
@@ 9 Theoretical Analysis
9.1 Consistency of Presence-Mask Normalization
@@
$S\;=\;\frac{w_{\text{r}}I_{\text{r}}\,s_{\text{ref}}+w_{\text{c}}I_{\text{c}}\,s_{\text{call}}+w_{\text{f}}I_{\text{f}}\,s_{\text{final}}}{W_{\text{act}}},\qquad W_{\text{act}}\;=\;w_{\text{r}}I_{\text{r}}+w_{\text{c}}I_{\text{c}}+w_{\text{f}}I_{\text{f}},$ (30)
@@
$S\;=\;\sum_{k\in\mathcal{A}}\alpha_{k}\,s_{k}\quad\text{with }s_{\text{r}}=s_{\text{ref}},\ s_{\text{c}}=s_{\text{call}},\ s_{\text{f}}=s_{\text{final}}.$ (32)
@@
9.2 Format Factor: Boundedness, Monotonicity, and EqualCalls Attenuation
@@
$P_{\text{total}}\;=\;P_{\text{miss}}\;+\;\beta_{\text{extra}}\,P_{\text{extra}}\;+\;\gamma_{\text{count}}\,P_{\text{count}},\qquad\beta_{\text{extra}},\gamma_{\text{count}}\geq 0,\quad P_{\bullet}\geq 0,$ (34)
@@
$r_{\mathrm{fmt}}\;=\;\begin{cases}r_{\text{reduce}},&\mathrm{EqualCalls}\!\big(C_{\text{calls}},\,G_{\text{calls}}\big),\\ 1,&\text{otherwise,}\end{cases}\qquad r_{\text{reduce}}\in(0,1].$ (35)
@@
$|\,\Delta F\,|\ \leq\ \lambda_{m}r_{\mathrm{fmt}}\,\bigl(\,|\Delta P_{\text{miss}}|+\beta_{\text{extra}}|\Delta P_{\text{extra}}|+\gamma_{\text{count}}|\Delta P_{\text{count}}|\,\bigr).$ (37)
9.3 Core Reward with Similarity Backoff: Signal and Variance Control
Let $R_{\text{core}}=S\cdot F$ as in (24). The total reward uses a backoff when $R_{\text{core}}$ is very small:
$R_{\text{total}}\;=\;\begin{cases}\mathrm{clip}_{[0,1]}(R_{\text{core}}),&R_{\text{core}}\geq\varepsilon,\\[4.0pt] \mathrm{clip}_{[0,1]}\!\left(w_{\text{b}}\cdot\mathrm{Sim}\!\big(\mathrm{concat}(C),\,\mathrm{concat}(G)\big)\right),&\text{otherwise,}\end{cases}$ (38)
with $w_{\text{b}}\in(0,1]$ and $\varepsilon>0$. Note $R_{\text{core}}\in[0,1]$ already, hence clipping is redundant but harmless and keeps the two branches notationally symmetric.

### v3

@@ 2.2 Reward Design
Let $\mathrm{EqualCalls}$ be the schema-strict equality on bags of calls. And let $E$ = $\mathrm{EqualCalls}(C_{\text{calls}},G_{\text{calls}})$ We use a reduction factor
$r\;=\;\begin{cases}r_{\text{reduce}},&\text{if }E,\\ 1,&\text{else,}\end{cases}\!,\quad r_{\text{reduce}}\in(0,1].$ (21)
Aggregate the penalty as
and define the instance-wise format factor. Let $P_{\text{fmt}}:=P_{\text{miss}}+P_{\text{extra}}+P_{\text{count}}$, $x]_{0}^{1}:=\min\{1,\max\{0,x\}\}$, and $F:=\mathrm{FormatFactor}(C,G)$. Then
$F=\begin{cases}1,&\text{if }P_{\text{fmt}}=0,\\ [1-\lambda_{m}P_{\text{total}}r]_{0}^{1},&\text{otherwise.}\end{cases}$ (23)
Here $\beta_{\text{extra}},\gamma_{\text{count}},\lambda_{m}\geq 0$ control the strength of the extra-part penalty, the count-mismatch penalty, and the overall scaling, respectively; $(w_{\text{ref}},w_{\text{calls}},w_{\text{final}})\geq 0$ are part weights.
Core reward and backoff.
The core reward is
$R_{\text{core}}\;=\;S\cdot F.$ (24)
Early in training, $S$ contains a binary component ($s_{\text{call}}\!\in\!\{0,1\}$) and $F$ applies hard penalties; small formatting or argument errors can drive $R_{\text{core}}$ close to zero. This yields sparse or unstable gradients and large variance across samples. To stabilize learning and provide a dense shaping signal when the exact-match objective is not yet achieved, we introduce a similarity backoff. Let $x]_{0}^{1}:=\max(0,\min(1,x))$ and $S_{\text{b}}:=\mathrm{Sim}\!\big(\mathrm{concat}(C),\,\mathrm{concat}(G)\big)$:
$R_{\text{total}}=\begin{cases}[\,R_{\text{core}}\,]_{0}^{1},&R_{\text{core}}\geq\varepsilon,\\ [\,w_{\text{b}}\,S_{\text{b}}\,]_{0}^{1},&\text{otherwise.}\end{cases}$ (25)
Here $w_{\text{b}}\in(0,1]$ and $\mathrm{concat}(\cdot)$ linearizes the messages.
2.3 RL for Tool-Reflection-Bench
We adopt a reinforcement-learning objective for tool calling that combines two complementary ideas: (i) DAPO-style decoupled clipping Yu et al. (2025): we use a decoupled clipping range with different lower/upper bounds $(\varepsilon_{\text{low}},\varepsilon_{\text{high}})$ and a clip-higher policy (a looser upper bound when $r>1$ for positive advantages), and we skip uninformative prompt groups whose rollouts carry negligible learning signal; (ii) GSPO-style sequence-level importance sampling Zheng et al. (2025): we compute the importance ratio at the sequence level and apply clipping at the same granularity as the sequence-level reward, which avoids the mismatch between token-wise importance sampling and sequence-level rewards and stabilizes optimization.
Objective.
$\mathcal{J}_{\text{RL}}(\theta)$ $=\mathbb{E}\!\left[\frac{1}{G}\sum_{i=1}^{G}\min\!\big(r_{i}(\theta),\,\bar{r}_{i}(\theta)\big)\,\hat{A}_{i}\right],$ (26)
$\bar{r}_{i}(\theta)$ $=\mathrm{clip}\!\big(r_{i}(\theta),1-\varepsilon_{\text{low}},1+\varepsilon_{\text{high}}\big).$
Here the expectation is over $(q,a)\sim\mathcal{D}$ and $\{o_{i}\}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q)$. We define $\mathrm{clip}(x,a,b)=\min\{b,\max\{a,x\}\}$ and typically take $\varepsilon_{\text{high}}>\varepsilon_{\text{low}}$ (“clip-higher”).
Prompt-group dynamic filtering.
DAPO skips prompt groups whose candidates provide almost no learning signal (e.g., all-correct or all-wrong). Concretely, let $\mu_{R}:=\mathrm{mean}(\{R_{j}\}_{j=1}^{G})$ and $\sigma_{R}:=\mathrm{std}(\{R_{j}\}_{j=1}^{G})$. We define batch-normalized advantages and a group-level acceptance set:
$\hat{A}_{i}$ $=\frac{R_{i}-\mu_{R}}{\sigma_{R}},$ (27)
$\mathcal{S}(q,a)$ $=\Big\{\,i\in\{1,\ldots,G\}:|\hat{A}_{i}|>\tau_{\text{adv}}\,\Big\}.$
We further require sufficient reward dispersion within the group:
$\mathrm{Var}\!\big(\{R_{i}\}_{i=1}^{G}\big)$ $>\tau_{\text{var}},$ (28)
$0<|\mathcal{S}(q,a)|$ $<G.$
Sequence-level importance ratio.
3 Experiments
@@ 3.1 Experiment Settings
We conduct training on our self-constructed Tool-Reflection-Bench. After human supervision and post-editing, we retained approximately 5k samples in JSONL format to ensure compatibility with RL training under the Swift Zhao et al. (2025b) framework.
@@
BFCL v3 Evaluation Metrics.
We follow the official BFCL v3 multi-turn evaluation. Each subset contains 200 multi-turn conversations, and we report conversation-level accuracy (pass@1): a conversation is correct only if it passes BFCL’s end-of-turn checks for all turns. BFCL executes the predicted tool calls and verifies correctness via both state-based (backend state) and response-based checks. We use the default termination rules (stop when no valid tool call is produced; force-terminate at 20 tool steps and mark incorrect). Multi-turn Overall is the unweighted average over the four subsets.
Tool-Reflection-Bench Evaluation Metrics.
To assess the model’s repair capability when tool calls fail, we used Tool-Reflection-Bench, with the evaluation metric being repair rate, Repair@n denotes that for the same data instance, if at least one out of n trials succeeds, the metric is recorded as 1; otherwise, it is 0.
@@
To verify the generalizability of Tool-Reflection-Bench and our training methodology, we conducted experiments using Llama3.1-8B Dubey et al. (2024), Qwen2.5-7B-Instruct Hui et al. (2024), and Qwen3-4B Yang et al. (2025) as base models.
@@
3.2 Experiment Results
Table 3: Ablation study on BFCL v3 multi-turn (conversation-level Pass@1 accuracy, %). We compare different RL variants on Qwen2.5-7B-Instruct-FC.
Base Model RL Method Base Miss_Func Miss_Param Long_Context Overall
Qwen2.5-7B-Instruct-FC / 16.50 11.00 9.00 7.50 11.00
DAPO 19.50 12.50 12.25 10.75 13.75
GSPO 20.25 11.50 11.75 9.50 13.25
[…]
Table 4: Ablation study on Tool-Reflection-Bench (Repair@n, %) for Llama-3.1-8B-Instruct. Prompt-only uses the inference prompt provided in the supplementary material (Appendix A.4).
Base Model Method Repair@1 Repair@3 Repair@5
Llama-3.1-8B-Instruct / 0.70 5.10 6.80
Prompt Only 1.50 9.30 12.60
SFT 3.20 12.40 16.90
[…]
3.2.1 Result on BFCL v3
Comparison with base models.
Table 1 reports BFCL v3 multi-turn accuracy for the base models and our post-training. Overall, our method consistently improves multi-turn tool-calling across all three backbones, with the largest gains on error patterns that require argument repair and retrieving missing information from long histories. On Llama-3.1-8B, Base increases from 5.0 to 9.5 (+95%) and Long_Context from 4.5 to 7.0 (+56%). On Qwen2.5-7B, we observe the largest improvement on Miss_Param (9.0 $\to$ 13.5, +50%), indicating stronger parameter correction under tool feedback. On Qwen3-4B, Multi-turn Overall increases from 16.25 to 20.75 (+28%), accompanied by a substantial Long_Context gain (+48%). In contrast, improvements on Miss_Func are smaller (e.g., 19.0 $\to$ 19.5 on Qwen3-4B), suggesting that selecting the correct tool/function remains harder than repairing arguments given a plausible tool choice, which matches the focus of our reflection-driven repair training.
3.2.2 Result on Tool-Reflection-Bench
Tool-Reflection-Bench evaluates repair under tool-call failures; the test split consists solely of perturbation-derived failure cases (rather than clean trajectories), thus measuring recovery rather than memorization. As shown in Table 2, open-source baselines are weak at one try (Repair@1 $\leq$ 9.6%) and improve only slightly with more attempts. Our post-training consistently boosts repair across all backbones: Llama-3.1-8B-Instruct improves from $0.7/5.1/6.8$ to 4.7/20.5/26.4, Qwen2.5-7B-Instruct from $2.4/6.1/8.0$ to 9.3/10.3/11.4, and Qwen3-4B-Instruct from $9.6/10.6/10.6$ to 14.9/18.5/19.5 (Repair@1/3/5). The larger gains at Repair@3/5 suggest more robust reflection-to-repair behavior under repeated attempts. Our fine-tuned models also outperform closed-source baselines (e.g., LongCat-Lite-8K-Chat Team et al. (2025), GPT-4o-mini OpenAI (2024b, a), GPT-4.1-mini OpenAI (2025)) across $n\!\in\!\{1,3,5\}$.
3.2.3 Ablation Studies
Analysis.
Table 3 shows that both DAPO and GSPO improve over the base model on BFCL v3 multi-turn, while our full method achieves the best overall performance across all four subsets. In particular, our gains are most pronounced on Miss_Param and Long_Context, consistent with our training signal that emphasizes reflection-driven argument repair and recovering missing information from long interaction histories. Table 4 further validates the necessity of learning-based post-training for repair: prompt-only self-correction provides a limited improvement, SFT yields larger gains, and RL delivers the strongest repair capability, especially under multiple trials (Repair@3/5), indicating that the reward-driven optimization better aligns the model with robust correction behavior beyond imitation alone.
@@ 4 Conclusion
5 Limitations
Generalization beyond designed failures.
Our training data is constructed via a set of perturbation-based disruptions that cover several common categories of tool-calling failures (e.g., missing/incorrect arguments and long-context dependencies). While our results on BFCL v3 multi-turn and Tool-Reflection-Bench demonstrate consistent improvements, these evaluations still represent a bounded set of failure modes. In real-world deployments, agents may encounter additional error types, such as ambiguous user intents, non-stationary tool APIs, partial tool outages, noisy tool outputs, or multi-tool plans with stronger temporal/causal dependencies. Extending perturbations and supervision to better approximate real failure distributions, as well as incorporating logs of naturally occurring failures, are promising directions.
Scaling to larger models.
We validate our method on 4B–8B class models and observe substantial gains. However, the behavior of much larger models (e.g., 70B+) may differ: stronger base capabilities can reduce headroom, and the optimal balance between supervised signals and RL shaping may change. Evaluating and adapting the training recipe for larger backbones—including whether lighter post-training (e.g., SFT-only or fewer RL steps) suffices—remains future work.
Appendix A Appendix
A.1 Use of LLMs
This work leveraged LLMs to verify the mathematical soundness and symbolic accuracy of a few formulas in Sec.A.8.
A.2 Related Works
A.2.1 Tool-augmented Large Language Models
Integrating external tools into large language models has become a key approach to enhancing their functionality, surpassing the simple task of text generation. Traditional LLMs are limited by static knowledge, constrained to the data they were trained on. However, tool-augmented models extend the capabilities of LLMs by enabling them to interact with external resources Zhang et al. (2024); Hao et al. (2025) (such as APIs Li et al. (2023), databases, and computational engines) through tool calls. This extension allows LLMs to access real-time data, perform external computations, and even interface with external hardware, making them more practical for solving complex real-world tasks that require dynamic information or specific external operations Chen et al. (2025a). ToolBench Qin et al. (2023) demonstrates the feasibility of integrating external tool calls into LLMs. Through such systems, LLMs can handle more specialized tasks. However, one major challenge of tool augmentation is how to effectively train LLMs to use these tools. Existing training methods, such as supervised fine-tuning and reinforcement learning, typically focus on optimizing single tool calls. This type of interaction often does not involve multi-turn tool calls or responses, which makes the limitations of current methods particularly apparent when errors occur during tool usage. In such cases, the model’s ability to recover from errors becomes crucial.
A.2.2 Self-Correction in LLMs
Self-correction in large language models refers to the model’s ability to diagnose its own errors and correct them based on previous actions Huang et al. (2023); Liu et al. (2024a). However, this area has not been fully explored. Existing self-correction techniques mostly rely on heuristic methods or unidirectional reasoning processes Renze and Guven (2024).
Self-Refine framework Madaan et al. (2023), which involves having LLMs provide an initial response, followed by a reflection process where the model identifies flaws and makes improvements. Specifically, the same LLM acts as both the responder and the evaluator: the model first generates an initial response, then self-reflects and iteratively revises the output. This approach has been shown to enhance the performance of LLMs in certain domains. However, subsequent studies Wu et al. (2024); Vladika et al. (2025) have found that relying solely on the model itself often fails to detect subtle errors. Some research Jiang et al. (2025); Zhao et al. (2025a) has introduced auxiliary verifiers (such as additional models or mechanisms Saveliev and Dendiuk (2024); Feng et al. (2025)) to help check the correctness of the initial response. This external self-checking assistance avoids unnecessary repeated revisions, improving efficiency and enhancing the model’s reasoning and verification capabilities. However, this approach remains highly sensitive to the specific phrasing of the prompts, with different prompt wordings leading to varying results Liu et al. (2024a).
Beyond prompt-only self-revision, recent work in tool-augmented agents improves reliability via interaction/execution feedback loops, e.g., unifying tool retrieval and calling via generation Wang et al. (2024), refining tool retrieval with iterative LLM feedback Xu et al. (2024), and trial-and-error frameworks that leverage tool interaction signals to iteratively improve tool understanding/documentation Qu et al. (2025).
However, even though these methods improve tool-use robustness, they primarily focus on leveraging online interaction signals (e.g., tool returns, success/failure) to drive behavior improvements. In contrast, our work targets self-correction itself as an explicit, trainable, and controllable capability: we formulate reflection-based error localization, diagnosis, and repair as a learning objective, and optimize it via supervised signals and reward shaping so that the model can reliably trigger and perform corrections during tool calling.
A.3 Cost–Benefit Analysis on Tool-Reflection-Bench
[…]
Llama-3.1-8B-Instruct / 0.70 5.10 6.80 /
DAPO 3.90 14.90 20.80 7h48m
GSPO 3.20 13.70 20.50 7h55m
[…]
Analysis.
[…]
A.4 Prompt for Prompt-only Self-correction
In this section, we provide a simplified prompt used for the Prompt Only baseline in our ablation studies. The prompt enforces a fixed output format (<reflect> + <call>) to encourage explicit diagnosis before proposing a corrected tool call.
How to run prompt-only self-correction
System Role. You are a tool-calling assistant. Goal. A previous tool call failed. Given the user goal, recent context, the failed tool call, and the tool error message, produce (1) a concise reflection diagnosing the failure, and (2) exactly one corrected tool call that is executable under the provided tool schema. Inputs. You will be given: [User Goal]: the task the assistant is trying to accomplish. [Tool API / Schema]: function names and argument schemas. [Context]: recent messages and tool results. [Failed Tool Call]: the tool call that triggered the error. [Tool Error / Feedback]: the tool response/error message. Output format (strict). Output must follow this exact format with no extra text: <reflect> - Error type: (wrong tool / missing tool / wrong arguments / missing arguments / wrong values / redundant call / long-context mismatch). - Evidence: cite the tool error and/or context that reveals the issue. - Fix: state the minimal change needed to repair the call. </reflect> <call>{"name": "...", "arguments": {...}}</call> Constraints. 1. Produce exactly one <reflect> block and exactly one <call> block. 2. The <call> must be valid JSON and executable under the provided schema. 3. Do not output <final>...</final> or any additional natural-language response. 4. Prefer minimal edits: only change what is necessary to fix the error.
User [User Goal] {USER_GOAL} [Tool API / Schema] {TOOL_SCHEMA_OR_DOCS} [Context (recent messages and tool results)] {CONTEXT} [Failed Tool Call] <call>{FAILED_CALL_JSON}</call> [Tool Error / Feedback] {ERROR_MESSAGE}
A.5 Prompt for Perturbation-based Disruptions
@@
A.5.1 Prompt for Call-Order Swap
@@
A.5.2 Prompt for Redundant Call
@@
A.5.3 Prompt for Missing Call
@@
A.5.4 Prompt for Argument Error
@@
A.6 Training Data Case Study
In this section, we extract one complete sample from each of the four perturbation modes for analysis. Due to space limitations, these four samples are provided in the supplementary material for reference, while here we only present a brief analysis of the data.
A.6.1 Case Study of Call-Order Swap
@@
A.6.2 Case Study of Redundant Call
@@
A.6.3 Case Study of Missing Call
@@
The tool schema exposes multiple functions with schema–strict parameters:
  • |check_plant_water_level(plant_location:string)|
  • |start_watering(plant_location:string, duration:number)|
  • |start_trimming(hedge_location:string)|
  • $\dots$
The user requests two primary actions in the backyard: (i) trim hedges and (ii) water all potted plants for about 10 minutes; afterwards ensure plants have enough water and dispose clippings
@@
  • |start_watering(plant_location="backyard", duration=10)|
  • |start_trimming(hedge_location="backyard")|
@@
A.7 Test Data Case Study
In this section, we present two representative test cases and their corresponding evaluation results as a case study, providing an intuitive demonstration of the effectiveness of our method and the model’s self-reflection capability for tool-call repair. Since the original cases are relatively long, we include their full content in the supplementary material for reference and provide only the analysis here.
A.7.1 Case I
@@ A.8 Theoretical Analysis
A.8.1 Consistency of Presence-Mask Normalization
@@
$W_{\text{act}}$ $=w_{\text{r}}I_{\text{r}}+w_{\text{c}}I_{\text{c}}+w_{\text{f}}I_{\text{f}},$ (30)
$S$ $=\frac{w_{\text{r}}I_{\text{r}}\,s_{\text{ref}}+w_{\text{c}}I_{\text{c}}\,s_{\text{call}}+w_{\text{f}}I_{\text{f}}\,s_{\text{final}}}{W_{\text{act}}}.$
@@
$\begin{split}S&=\sum_{k\in\mathcal{A}}\alpha_{k}\,s_{k},\\ &\text{where }s_{\text{r}}=s_{\text{ref}},\;s_{\text{c}}=s_{\text{call}},\;s_{\text{f}}=s_{\text{final}}.\end{split}$ (32)
@@
A.8.2 Format Factor: Boundedness, Monotonicity, and EqualCalls Attenuation
@@
$\begin{split}P_{\text{total}}&=P_{\text{miss}}+\beta_{\text{extra}}\,P_{\text{extra}}+\gamma_{\text{count}}\,P_{\text{count}},\\ &\text{where }\beta_{\text{extra}},\gamma_{\text{count}}\geq 0\text{ and }P_{\bullet}\geq 0.\end{split}$ (34)
@@
$\begin{split}\text{Let }E&:=\mathrm{EqualCalls}\!\big(C_{\text{calls}},\,G_{\text{calls}}\big).\\[-1.0pt] r_{\mathrm{fmt}}&=\begin{cases}r_{\text{reduce}},&E,\\ 1,&\text{otherwise,}\end{cases}\qquad r_{\text{reduce}}\in(0,1].\end{split}$ (35)
@@
$\begin{split}|\,\Delta F\,|\ \leq\ \lambda_{m}r_{\mathrm{fmt}}\Big(|\Delta P_{\text{miss}}|+\beta_{\text{extra}}|\Delta P_{\text{extra}}|\\ \qquad\qquad\qquad+\gamma_{\text{count}}|\Delta P_{\text{count}}|\Big).\end{split}$ (37)
A.8.3 Core Reward with Similarity Backoff: Signal and Variance Control
Let $R_{\text{core}}=S\cdot F$ as in (24). To stabilize learning when $R_{\text{core}}$ is very small, we introduce a similarity backoff. Define $x]_{0}^{1}:=\min\{1,\max\{0,x\}\}$ and $S_{\text{b}}:=\mathrm{Sim}\!\big(\mathrm{concat}(C),\,\mathrm{concat}(G)\big)$:
$R_{\text{total}}=\begin{cases}[\,R_{\text{core}}\,]_{0}^{1},&R_{\text{core}}\geq\varepsilon,\\ [\,w_{\text{b}}\,S_{\text{b}}\,]_{0}^{1},&\text{otherwise.}\end{cases}$ (38)
Here $w_{\text{b}}\in(0,1]$ and $\varepsilon>0$. Note that $R_{\text{core}}\in[0,1]$ already, so clipping is redundant but harmless, and it keeps the two branches notationally symmetric.